# Constrained Diffusion for Protein Design with Hard Structural Constraints

### v1

@@ 1 Introduction
These limitations highlight the broader challenge of designing structured objects under strict feasibility. Many existing approaches augment diffusion models through constraint guidance (Ho et al., 2020; Ho & Salimans, 2022). For example, Gruver et al. (2023) inject soft constraints into discrete sequence prediction for protein design through gradient-based guidance; however, while guidance approaches result in increased feasibility rates, they fail to consistently provide constraint adherent outputs. Others adopt post-processing optimizations which more rigorously target the constraint set. However, these methods rely on either simplifications of the highly nonconvex constraint set (e.g., matching an existing ligand template) or result in samples falling outside the data manifold (Bergues et al., 2025; Christopher et al., 2024). A more effective approach is to inject constraint into the generative process, e.g., by projecting intermediate states back to the feasible set (Christopher et al., 2024). Yet projecting noisy states early in the sampling process has been show to disrupt the diffusion trajectory, potentially biasing samples (Blanke et al., 2025).
@@ 3 Challenges with Constrained Diffusion
    As observed by Blanke et al. (2025), projections on intermediate states introduces statistical biases. This arises as intermediate samples concentrate near constraint boundaries. This issue has also been reported for soft guidance schemes, where increased weight on the guidance terms tends to disrupt the diffusion trajectory and degrading performance (Dhariwal & Nichol, 2021; Nichol & Dhariwal, 2021; Ho & Salimans, 2022).
@@
    Additionally, intermediate feasibility requires projecting highly noisy states onto complex, nonconvex constraints, which can result in solutions trapped in local minima (Pardalos & Vavasis, 1991; Boyd & Vandenberghe, 2004), disrupting the diffusion performance while simultaneously limiting the efficacy of feasibility updates.
@@ 4 Reverse Diffusion as Proximal Optimization
A natural first choice would be to take $g$ as the indicator function of the feasible set. However, this would require exact projections onto a potentially nonconvex set, which can be ill-defined when $\hat{\bm{x}}_{0}^{t}$ lies far from $\mathcal{C}$. To avoid instability, the hard indicator is replaced by its Moreau envelope (Boyd & Vandenberghe, 2004), yielding the smooth penalty
@@ 5 Decoupling Global Topology from Local Geometry via ADMM
The constraint set $\mathcal{C}$ captures a strong coupling between local stereochemical variables and global variables governing topology and long-range residue interactions. Because residues that are far apart in sequence may lie adjacent in the folded structure, enforcing global constraints thus necessitates coordinated updates that can significantly impact nearby stereochemisty. For instance, we observe that applying covalent bond constraints on a $\beta$-strand can cause the associated residues to shift substantially, degrading the fidelity of this local geometry (Budyak et al., 2024). These interdependencies make the proximal step computationally complex. However, the presence of these separable local and global constraints confers structure to the problem and thus presents an opportunity to exploit it, enabling the use of decomposition approaches.
Consider that a feasible point can be equivalently represented as $\bm{x}\in\mathcal{C}_{\mathrm{local}}\cap\mathcal{C}_{\mathrm{global}}$. The local constraints $\mathcal{C}_{\mathrm{local}}$ capture properties that are applicable in all backbone design tasks (e.g., adherence to stereochemical bond lengths and angles between consecutive atoms and residues). The global constraints $\mathcal{C}_{\mathrm{global}}$ are problem-specific functionals: for example, in our first experiment, this constraint set defines bond lengths and angles between specific non-neighboring residues, characterizing covalent bonds which are necessary for protein-ligand pocket design.
@@
These updates can, thus, be interpreted as applying ADMM to the consensus problem, where the dual variable $\bm{u}$ carries forward residuals between local and global feasibility corrections. In practice, this is implemented by minimizing the associated augmented Lagrangian. For clarity, we present only the proximal form here but detail the explicit augmentented Lagrangian in Appendix E.
@@ 6 Theoretical Analysis
Now, we show that the samples generated by the proposed stochastic proximal method come with feasibility guarantees. In the following we assume that the constraint set $\mathcal{C}$ is prox-regular and defer all proofs in Appendix F.
@@ 7 Experiments
    Constraint Satisfaction: The percentage of samples satisfying all domain-specific constraints. For Section 7.1, we verify presence of the described covalent bonds through DSSP (Kabsch & Sander, 1983), while Section 7.2 computes constraints via Cartesian coordinate bounds.
@@
7.1 PDZ Domain: Covalent Bonds
@@
Results. Table 1 provides results on the PDZ benchmark comparing our constrained diffusion approach (visualized in Figure 2 (a)) to the baseline methods. Notably, across nearly one hundred thousand samples generated for the three baselines, not one sample perfectly satisfied the bonding distance and angle constraints. The baselines frequently generate incorrect secondary structures, as illustrated in Figure 2 (b), making it implausible that generations will effectively bind with the peptide ligand. While recentering and CGD perform well in terms of the local geometric requirements captured by the structure realism measurement, they are unable to cope with the global requirement of covalent bonding between the PDZ backbone continuation and the peptide ligand, as reflect by the constraint satisfaction rates. While constraint-guided diffusion satisfies the bond distance constraints for some generations, it is never able to generate residues which appropriately meet the angle requirements. Importantly, the baselines yield no usable generations for any of the 31 structures in the benchmark. In contrast, our method achieves state-of-the-art results, generating usable structures in 21.0% of total generations (and up to 83.0% for well-posed ligands), markedly outperforming the existing baselines. In addition to observing perfect constraint satisfaction, we outperform all methods substantially in radius of gyration and diversity metrics. These results highlight the unique ability of our approach to handle both local stereochemical properties while enforcing global functional constraints, providing a vastly more viable approach to protein engineering under specific property requirements and design constraints.
@@ 7.2 Molecule Encapsulation: Vacancy Constraints
Results. Figure 3 provides a visualization of representative samples from each baseline. Standard diffusion and recentering of mass guidance perform similarly in this domain, yielding realistic structures but failing to satisfy the functional requirements in any out of 1000 samples generated for each. We observe slightly different failure modes: standard diffusion generally violates the box constraint, while recentering more often violates the vacancy constraint. The recentering is effective at keeping the structures inside the box, but it cannot capture the exclusion zone. Table 2 reports much stronger performance for constraint-guided diffusion, which generates feasible samples 53.7% of the time, resulting in 52.6% usable samples. However, it is worth noting that the qualitative performance suffers, as the radius of gyration is much higher than other baselines. This is often indicative of structures which contain unfolded conformations, ultimately undermining structural stability and realism. In comparison, our method reports perfect constraint satisfaction, producing 4.8 $\times$ more usable samples as the nearest baseline, with an impressive 97.8% success rate. Furthermore, it maintains radii of gyration comparable to standard diffusion, indicating the generated samples combine structural plausibility, compactness, and fold coverage.
@@
Standard Recenter CGD
Constraint Satisfaction (%) $\uparrow$] 0.0 0.0 21.6 100.0
Structure Realism (%) $\uparrow$] (100.0) (100.0) 96.1 97.8
Usable Percentage (%) $\uparrow$] 0.0 0.0 20.5 97.8
Radius of Gyration (Å) $\downarrow$] (15.2) (14.3) 23.9 14.8
Diversity (%) $\uparrow$] N/A N/A 20.5 97.8
@@ 8 Related Work
While existing de novo protein structure design models produce plausible generations, as we have shown in this paper, sampled protein backbones frequently violate inter-atomic bond lengths, angles, or chain closure requirements, often necessitating the generation of tens of thousands of candidates to obtain a handful of viable designs (e.g., Watson et al. 2023; Sappington et al. 2024). Although backbone generators such as RFdiffusion (Watson et al., 2023; Ahern et al., 2025) provide major advances in functional conditioning, outputs still require post hoc filtering to ensure stereochemical correctness, and current pipelines continue to rely heavily on rejection sampling.
@@
Inference-time approaches have emerged as a strong alternative, enforcing per-sample compliance and removing the need for model retraining. Diffusion guidance methods were first introduced for soft constraint imposition but are fundamentally limited, offering only probabilistic bias rather than guaranteed adherence to the constraint set (Ho & Salimans, 2022). While these techniques have improved performance for protein backbone generation, with models such as Chroma (Ingraham et al., 2023) leveraging conditioning on specific substructures, as we have shown in our experiments, they are often ineffective in providing consistent constraint satisfaction. Overcoming these limitations requires generative models which can effectively integrate these constraints into the design process, as presented in this work.
@@ Acknowledgments
This research is partially supported by NSF awards 2533631, 2401285, 2334936, and 2226816, by DARPA under Contract No. #HR0011252E005, and by the National Institutes of Health, National Institute of General Medical Sciences, under Award Number T32 GM135141. The authors acknowledge the Research Computing at the University of Virginia. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of NSF, DARPA, or the National Institutes of Health.
@@ Appendix A Limitations and Future Work
In scientific domains such as protein engineering, exact constraint satisfaction is essential for both physical realism (local constraints) and functional utility (global constraints). However, it is valuable to acknowledge several important trade-offs associated with the adoption of constrained methods. First, enforcing feasibility inherently introduces increased computational complexity. Sampling from $p_{\mathcal{C}}$ is a fundamentally harder task than sampling from $p_{\mathrm{data}}$, especially under the assumption of hard constraint satisfaction, as reflected in reported runtimes (Appendix D). While runtime is worth considering, especially as the method scales, it is important to note that when sampling from $p_{\mathcal{C}}$, our approach offers the most efficient runtime, providing much better performance than the current rejection sampling paradigm. Although not applicable in our experimental settings, if sampling from $p_{\mathrm{data}}$ is sufficient, it is not computationally tractable to adopt constrained diffusion methods.
@@ Appendix B Experimental Details
B.1 PDZ Domain: Covalent Bonds
@@
where $c_{i}$ denotes the score of constraint violation $i$. Earlier in the diffusion process, $\beta=30$ remains high, keeping the resampling fairly stochastic, but we lower this $\beta=1$ as $t\rightarrow 0$ to improve selection for the final state and increase the likelihood of covalent bonds forming.
@@ Appendix C Dataset Construction
Appendix D Runtime

### v2

@@ 1 Introduction
These limitations highlight the broader challenge of designing structured objects under strict feasibility. Many existing approaches augment diffusion models through constraint guidance (Ho et al., 2020; Ho and Salimans, 2022). For example, Gruver et al. (2023) inject soft constraints into discrete sequence prediction for protein design through gradient-based guidance; however, while guidance approaches result in increased feasibility rates, they fail to consistently provide constraint adherent outputs. Others adopt post-processing optimizations which more rigorously target the constraint set. However, these methods rely on either simplifications of the highly nonconvex constraint set (e.g., matching an existing ligand template) or result in samples falling outside the data manifold (Bergues et al., 2025; Christopher et al., 2024). A more effective approach is to inject constraint into the generative process, e.g., by projecting intermediate states back to the feasible set (Christopher et al., 2024). Yet projecting noisy states early in the sampling process has been show to disrupt the diffusion trajectory, potentially biasing samples (Blanke et al., 2025).
@@ 3 Challenges with Constrained Diffusion
    As observed by Blanke et al. (2025), projections on intermediate states introduces statistical biases. This arises as intermediate samples concentrate near constraint boundaries. This issue has also been reported for soft guidance schemes, where increased weight on the guidance terms tends to disrupt the diffusion trajectory and degrading performance (Dhariwal and Nichol, 2021; Nichol and Dhariwal, 2021; Ho and Salimans, 2022).
@@
    Additionally, intermediate feasibility requires projecting highly noisy states onto complex, nonconvex constraints, which can result in solutions trapped in local minima (Pardalos and Vavasis, 1991; Boyd and Vandenberghe, 2004), disrupting the diffusion performance while simultaneously limiting the efficacy of feasibility updates.
@@ 4 Reverse Diffusion as Proximal Optimization
A natural first choice would be to take $g$ as the indicator function of the feasible set. However, this would require exact projections onto a potentially nonconvex set, which can be ill-defined when $\hat{\bm{x}}_{0}^{t}$ lies far from $\mathcal{C}$. To avoid instability, the hard indicator is replaced by its Moreau envelope (Boyd and Vandenberghe, 2004), yielding the smooth penalty
@@ 5 Decoupling Global Topology from Local Geometry via ADMM
The constraint set $\mathcal{C}$ captures a strong coupling between local stereochemical variables and global variables governing topology and long-range residue interactions. Because residues that are far apart in sequence may lie adjacent in the folded structure, enforcing global constraints thus necessitates coordinated updates that can significantly impact nearby stereochemisty. For instance, we observe that applying non-covalent bond constraints on a $\beta$-strand can cause the associated residues to shift substantially, degrading the fidelity of this local geometry (Budyak et al., 2024). These interdependencies make the proximal step computationally complex. However, the presence of these separable local and global constraints confers structure to the problem and thus presents an opportunity to exploit it, enabling the use of decomposition approaches.
Consider that a feasible point can be equivalently represented as $\bm{x}\in\mathcal{C}_{\mathrm{local}}\cap\mathcal{C}_{\mathrm{global}}$. The local constraints $\mathcal{C}_{\mathrm{local}}$ capture properties that are applicable in all backbone design tasks (e.g., adherence to stereochemical bond lengths and angles between consecutive atoms and residues). The global constraints $\mathcal{C}_{\mathrm{global}}$ are problem-specific functionals: for example, in our first experiment, this constraint set defines bond lengths and angles between specific non-neighboring residues, characterizing non-covalent bonds which are necessary for protein-ligand pocket design.
@@
These updates can, thus, be interpreted as applying ADMM to the consensus problem, where the dual variable $\bm{u}$ carries forward residuals between local and global feasibility corrections. In practice, this is implemented by minimizing the associated augmented Lagrangian. For clarity, we present only the proximal form here but detail the explicit augmentented Lagrangian in Appendix G.
@@ 6 Theoretical Analysis
Now, we show that the samples generated by the proposed stochastic proximal method come with feasibility guarantees. In the following we assume that the constraint set $\mathcal{C}$ is prox-regular and defer all proofs in Appendix H.
@@ 7 Experiments
    Constraint Satisfaction: The percentage of samples satisfying all domain-specific constraints. For Section 7.1, we verify presence of the described hydrogen-bond pairing of the backbone through DSSP (Kabsch and Sander, 1983), while Section 7.2 computes constraints via Cartesian coordinate bounds.
@@
7.1 PDZ Domain: Non-covalent Bonds
@@
Results. Table 1 provides results on the PDZ benchmark comparing our constrained diffusion approach (visualized in Figure 2 (a)) to the baseline methods. Notably, across nearly one hundred thousand samples generated for the three baselines, not one sample perfectly satisfied the bonding distance and angle constraints. The baselines frequently generate incorrect secondary structures, as illustrated in Figure 2 (b), making it implausible that generations will effectively bind with the peptide ligand. While recentering and CGD perform well in terms of the local geometric requirements captured by the structure realism measurement, they are unable to cope with the global requirement of non-covalent bonding between the PDZ backbone continuation and the peptide ligand, as reflect by the constraint satisfaction rates. While constraint-guided diffusion satisfies the bond distance constraints for some generations, it is never able to generate residues which appropriately meet the angle requirements. Importantly, the baselines yield no usable generations for any of the 31 structures in the benchmark. In contrast, our method achieves state-of-the-art results, generating usable structures in 21.0% of total generations (and up to 83.0% for well-posed ligands), markedly outperforming the existing baselines. In addition to observing perfect constraint satisfaction, we outperform all methods substantially in radius of gyration and diversity metrics. These results highlight the unique ability of our approach to handle both local stereochemical properties while enforcing global functional constraints, providing a vastly more viable approach to protein engineering under specific property requirements and design constraints.
@@ 7.2 Molecule Encapsulation: Vacancy Constraints
[…]
@@
Standard Recenter CGD Recenter + CGD
Constraint Satisfaction (%) $\uparrow$] 0.0 0.0 21.6 27.4 100.0
Structure Realism (%) $\uparrow$] (100.0) (100.0) 96.1 93.8 97.8
Usable Percentage (%) $\uparrow$] 0.0 0.0 20.5 24.2 97.8
Radius of Gyration (Å) $\downarrow$] (15.2) (14.3) 23.9 26.6 14.8
Diversity (%) $\uparrow$] N/A N/A 20.5 24.2 97.8
@@ 8 Related Work
While existing de novo protein structure design models produce plausible generations, as we have shown in this paper, sampled protein backbones frequently violate inter-atomic bond lengths, angles, or chain closure requirements, often necessitating the generation of tens of thousands of candidates to obtain a handful of viable designs (e.g., Watson et al. 2023; Sappington et al. 2024). Although backbone generators such as RFDiffusion (Watson et al., 2023; Ahern et al., 2025) provide major advances in functional conditioning, outputs still require post hoc filtering to ensure stereochemical correctness, and current pipelines continue to rely heavily on rejection sampling. Recent models such as Genie 2 and OriginFlow further illustrate the growing interest in diffusion and flow-based approaches for backbone generation (Lin et al., 2024; Yan et al., 2025); however, these methods exhibit similar limitations, with generated backbones often violating global constraints and requiring substantial downstream filtering.
@@
Inference-time approaches have emerged as a strong alternative, enforcing per-sample compliance and removing the need for model retraining. Diffusion guidance methods were first introduced for soft constraint imposition but are fundamentally limited, offering only probabilistic bias rather than guaranteed adherence to the constraint set (Ho and Salimans, 2022). While these techniques have improved performance for protein backbone generation, with models such as Chroma (Ingraham et al., 2023) leveraging conditioning on specific substructures, as we have shown in our experiments, they are often ineffective in providing consistent constraint satisfaction. Overcoming these limitations requires generative models which can effectively integrate these constraints into the design process, as presented in this work.
@@ 9 Conclusion
Ethics Statement
This work develops methods for constrained generative modeling in protein design. Our method improves feasibility in backbone design tasks, which holds significant potential to accelerate existing protein engineering pipelines. To mitigate risks of potential misuse, this paper restricts evaluation to safe, publicly available structural benchmarks, curated from the Protein Data Bank. All implementations follow standard open science practices, ensuring safe and transparent research
Reproducibility Statement
The code and benchmark are released in the supplementary material along with instructions to guide the reproduction of the results presented in this paper. Methodological details are described extensively in the paper and accompanying appendix. Section 7, Appendix B, Appendix F describe the specific constraints, hyperparameters, models, and hardware used for evaluation pipelines.
@@ Acknowledgments
This research is partially supported by NSF awards 2533631, 2401285, 2334936, and 2226816, by DARPA under Contract No. #HR0011252E005, by the National Institutes of Health, National Institute of General Medical Sciences, under Award Number T32 GM135141, and Sargassum BioRefinery (SaBRe) Center, a project of Schmidt Sciences’ Virtual Institute of Feedstocks of the Future (VIFF), with support from the Foundation for Food & Agriculture Research. The authors acknowledge the Research Computing at the University of Virginia. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of NSF, DARPA, or the National Institutes of Health.
@@ Appendix A Limitations and Future Work
In scientific domains such as protein engineering, exact constraint satisfaction is essential for both physical realism (local constraints) and functional utility (global constraints). However, it is valuable to acknowledge several important trade-offs associated with the adoption of constrained methods. First, enforcing feasibility inherently introduces increased computational complexity. Sampling from $p_{\mathcal{C}}$ is a fundamentally harder task than sampling from $p_{\mathrm{data}}$, especially under the assumption of hard constraint satisfaction, as reflected in reported runtimes (Appendix F). While runtime is worth considering, especially as the method scales, it is important to note that when sampling from $p_{\mathcal{C}}$, our approach offers the most efficient runtime, providing much better performance than the current rejection sampling paradigm. Although not applicable in our experimental settings, if sampling from $p_{\mathrm{data}}$ is sufficient, it is not computationally tractable to adopt constrained diffusion methods.
@@ Appendix B Experimental Details
B.1 PDZ Domain: Non-covalent Bonds
@@
where $c_{i}$ denotes the score of constraint violation $i$. Earlier in the diffusion process, $\beta=30$ remains high, keeping the resampling fairly stochastic, but we lower this $\beta=1$ as $t\rightarrow 0$ to improve selection for the final state and increase the likelihood of non-covalent bonds forming.
@@ B.2 Molecule Encapsulation: Vacancy Constraints
Additional Results.
To supplement the results reported in the paper, we include additional analysis of the secondary structure preservation rates for samples generated by each baseline and by our method. This metric quantifies the percentage of residues assigned to a canonical secondary structure state in the relaxed models. Higher values indicate improved preservation of fold-level features.
RFDiffusion Ours
Standard Recenter CGD
Secondary Structure (%) 80.8 82.5 83.5 85.0
Table 4: Secondary structure preservation across design methods.
In addition to the primary metrics reported in the main text, we assess the structural plausibility of generated designs using a secondary structure preservation metric. This measure is assessed using PyRosetta to determine the percentage of residues which fall in helix, sheet, or turn conformations in the relaxed structures (Chaudhury et al., 2010). As shown in Table 4, our method achieves the highest secondary structure frequency, suggesting that the model maintains global fold characteristics more reliably than the baselines.
@@ Appendix C Dataset Construction
Appendix D Comparison to Projected Diffusion Models
As discussed in Section 3, prior work on constrained generative sampling has proposed Projected Diffusion Models (Christopher et al., 2024). These models augment the sampling process by imposing a projection onto the feasible set after each diffusion step. While theoretical guarantees have been derived for convex constraint sets, and empirical observations have illustrated that these properties hold for certain nonconvex sets, the paper has argued critical limitations for step-then-project approaches when applied to challenging settings as explored in this work. This section presents empirical justification of these aforementioned claims.
Projected Diffusion Model Ours
Local + Global Global
Constraint Satisfaction (%) 0.0 (Failed to Converge) 100.0 100.0
Structure Realism (%) N/A 0.0 97.8
Usable Percentage (%) 0.0 0.0 97.8
Radius of Gyration (Å) N/A (13.1) 14.8
Diversity (%) N/A N/A 97.8
Table 5: Comparison to Projected Diffusion Models on the Molecule Encapsulation setting.
The results in Table 5 reaffirm the limitations of integrating intermediate projection steps when working with high-dimensional data subject to complex nonconvex constraints. While in this setting it was possible to exactly satisfy the global constraints, the resulting structures are unusable, with high concentrations of atoms placed on the edges of the box and cone boundaries. When also enforcing the local constraints explicitly, the projection operator never converged on any samples or timesteps.
Appendix E Hyperparameter Sensitivity Analysis
To assess the robustness of our method to hyperparameter choices, we performed two ablation studies: (1) varying constraint weight schedule $\lambda$, and (2) varying the number of ADMM iterations per diffusion step. All experiments were conducted on the PDZ domain design target 6ubh, a well-posed example from our benchmark set suited to constrained generation.
$\lambda=\tfrac{10}{t}$ $\lambda=\tfrac{100}{t}$ $\lambda=\tfrac{1000}{t}$ $\lambda=\inf$
Constraint Satisfaction (%) 100.0 100.0 100.0 100.0
Structure Realism (%) 80.0 100.0 100.0 90.0
Usable Percentage (%) 80.0 100.0 100.0 90.0
Radius of Gyration (Å) 10.58 10.53 10.56 10.34
Diversity (%) 80.0 100.0 100.0 90.0
Table 6: Ablation over constraint-weight schedules $\lambda$ on PDZ domain sample 6ubh.
As illustrated by Table 6, increasing the initial weight of $\lambda_{T}$ yields improvements in structural realism and usable sample percentage. This trend aligns with the interpretation that stricter constraint enforcement at early timesteps guides the diffusion trajectory toward the feasible manifold, improving geometric quality. However, increasing this too strictly, results in over constraining early states, analogous to placing an overly large weight on the diffusion-guidance terms as discussed in Section 3.
ADMM = 1 ADMM = 2 ADMM = 3
Constraint Satisfaction (%) 100.0 100.0 100.0
Structure Realism (%) 80.0 70.0 40.7
Usable Percentage (%) 80.0 70.0 40.7
Radius of Gyration (Å) 10.58 10.72 11.24
Diversity (%) 80.0 70.0 40.7
Table 7: Ablation on the number of ADMM iterations per diffusion step.
The next analysis centers on determining the optimal number of ADMM iterations per diffuson step. As described in Section 5, a single ADMM sweep per diffusion step is sufficient in practice. In fact, the ablation suggests that multiple iterations degrade quality. The failure case that emerges over additional ADMM iterations is actually always tied to the formation of a $\beta$-sheet across the binding sites; additional iterations do not result in local stereochemical properties breaking, but the analysis suggests that they do lower the liklihood of secondary structure conformations. This is because a higher wieght is placed on the ADMM optimization than the diffusion updates, leading to feasible generations but not ones which fall closely on the data manifold. Furthermore, we note that additional ADMM steps also become impractical because adding these iterations scales the runtime nearly linearly.
Appendix F Runtime